Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

198 700+ problem question 16

Expert replies
by Psyker » Thu Jun 17, 2010 9:37 am
"A basket has 5 apples, one is spoiled. If Henry picked 2 apples simultaneously and at random, what is the probability that the 2 selected apples will include the spoiled one.

a) 1/5
b) 2/10
c) 2/5
d) 1/2
e) 3/5

Hi guys,

I'm kinda stuck on the reasoning for the answer being "C". Intuitively I thought 2/5 because in my reasoning, picking 2 simultaneously doesn't change the fact that 1/5 apples is always spoiled, and the remaining apple that can be picked to make up the 2 is 1/5?

Not sure if my reasoning makes sense. Could some please help to provide the explanation for this question? Thanks much in advance!
Join the discussion
Source: — Problem Solving |

by asamaverick » Thu Jun 17, 2010 9:55 am
Here is how I would approach this.

Picking 2 apples from the basket containing 5 : You can select this in 5C2 = 10 ways.
Out of these 10 ways, there will be 4 instances in which the spoiled one is selected.
If you denote the apples as G1, G2, G3, G3 & S. The four combinations of picking spoiled one are SG1, SG2, SG3 & SG4. (Remember order does not matter since we are picking them simultaneously SG1 = G1S).

So the probability = 4/10 = 2/5.
Join the discussion

by jube » Thu Jun 17, 2010 10:08 am
Another approach:

Probability that one of the apples will be spoiled = 1- Probability that neither of the apples are spolit

Probability that neither of the apples are spoilt: (4/5)(3/4)=3/5

Answer: 1-(3/5)=2/5
Join the discussion

by kvcpk » Thu Jun 17, 2010 10:25 am
I am not sure if my approach is rite wor wrong.. I want an expert to correct me..

The question is same as picking 1 apple each time for 2 times.. but with an OR condition.

So the first time when i pick, probability to pick the spoilt apple is 1/5.
second time to pick it is also 1/5 [from the same apple set again]

we know that one of these two should occur. So 1/5 + 1/5 = 2/5

Please let me know if i am wrong!!
Join the discussion

by indiantiger » Thu Jun 17, 2010 1:26 pm
the probability of getting a spoiled apple is 1/5

we can have total two cases :

Case 1 first one is spoil and second is fine.

Case 2 first one is fine and second one is spoil

so total cases * probability of the event

2 * 1/5 = 2/5

Please correct me if I am wrong.
"Single Malt is better than Blended"
Join the discussion

by sk818020 » Thu Jun 17, 2010 3:10 pm
I solved it rather quickly by:

Probability you the first one you draw is poisoned, a = 1/5
Probability the second is not poisoned, b = 4/4

1/5*4/4=4/20

Probability the first you draw is not poisoned, c = 4/5
Probability the second drawn is poisoned, d = 1/4

4/5*1/4=4/20

4/20+4/20=8/20=2/5

Or using statistical notation;

[P(a)*P(b)]+[P(c)*P(d)]=4/20+4/20=2/5

Hope that helps.

Thanks,

Jared
Join the discussion

by artistocrat » Thu Jun 17, 2010 5:37 pm
I really like some of the methods for solving this problem that are described here. Here is another way to look at the problem.

So you have five apples: 1,2,3,4,5 (let's say 1 is rotten)

You have 10 pairs (taking out 2 "simultaneously"):

12 23 34 45
13 24 35
14 25
15

Notice that 1 is found in 4 pairs of 10.

4/10=2/5
Join the discussion

by artistocrat » Thu Jun 17, 2010 5:39 pm
I really like some of the methods for solving this problem that are described here. Here is another way to look at the problem.

So you have five apples: 1,2,3,4,5 (let's say 1 is rotten)

You have 10 pairs (taking out 2 "simultaneously"):

12 23 34 45
13 24 35
14 25
15

Notice that 1 is found in 4 pairs of 10.

4/10=2/5
Join the discussion

by Psyker » Thu Jun 17, 2010 6:22 pm
Thanks much to everyone for their help! :) really useful to see the different approaches.
Join the discussion

by kvcpk » Thu Jun 17, 2010 9:58 pm
kvcpk wrote:I am not sure if my approach is rite wor wrong.. I want an expert to correct me..

The question is same as picking 1 apple each time for 2 times.. but with an OR condition.

So the first time when i pick, probability to pick the spoilt apple is 1/5.
second time to pick it is also 1/5 [from the same apple set again]

we know that one of these two should occur. So 1/5 + 1/5 = 2/5

Please let me know if i am wrong!!
Can some expert confirm if my approach above is correct?
Join the discussion